How much mass defect is required to release 5.5 x 1020 J of energy?

A reaction takes place with a mass defect of 0.030kg. How much energy does this correspond to?

A reaction of 2.5kg of fissionable material takes place. Assuming a 0.10% mass defect, how much energy is released?
Answer #1

1) To Convert joule into eV one has to consider the factor 6.2 * 1018

1 J = 6.2 * 1018 eV

Thus 5.5 *1020 J = 6.2* 5.5 * 1018+20= 34.1 * 1038 eV = 34.1 * 1032 MeV

According to mass defect relation:

E= MeV

34.1 * 1032 MeV = 931.5 *

= 3.6 * 1030 amu

2) 0.030 kg = 1.80 * 1025 amu.

E= MeV

= 931.5 * 1.8* 1025 MeV

= 1676.7 * 1025 MeV.

3) 0.10% of 2.5kg = 0.25kg ( mass defect)

0.25 kg = 1.5 * 1026 amu

E= MeV

= 931.5 * 1.5 * 1026 MeV

=1397.25 * 1026 MeV